With an applied force of 225 N, the car will accelerate and move from rest travelling 25 m in 19.5 s. what is the mass of the car
1
Expert's answer
2019-11-04T10:43:29-0500

The acceleration is


"a=\frac{2h}{t^2}=\frac{2(25)}{19.5^2}=0.1315\frac{m}{s^2}"

The mass of the car:


"m=\frac{F}{a}=\frac{225}{0.1315}=1710\ kg"